If i have a trapezoid and the top is 12 the bottom is 30 what is the middle line
1
Expert's answer
2013-01-18T08:50:24-0500
The middle line of a trapezoid is equal to the ariphmeticmean of the top and the bottom sides:

m = (t+b)/2 =(12+30)/2 = 21